OpenStudy (anonymous):

find the exact value of each expression. tan(sin^(-1)(x/2))

OpenStudy (freckles):

let u=sin^(-1)(x/2) so sin(u)=x/2 draw right triangle

OpenStudy (freckles):

you will need to find the adjacent side using Pythagorean theorem

OpenStudy (freckles):

then you can find tan(u) and you will be done

OpenStudy (anonymous):

Ok thanks very much. I understand now

OpenStudy (freckles):

did you want me to check your answer?

OpenStudy (anonymous):

is it x/(4-x^2)?

OpenStudy (freckles):

almost

OpenStudy (freckles):

did you put a sqrt( ) on the bottom of all that one junk

OpenStudy (freckles):

x/sqrt(4-x^2)

OpenStudy (anonymous):

so if I put the sqrt the answer is correct

OpenStudy (freckles):